[Halld] [New Logentry] RC summary for past 24 hours

aasatury at jlab.org aasatury at jlab.org
Sat Oct 22 11:05:01 EDT 2022


Logentry Text:
--
Over the past 24 hours we collected ~3.43B total triggers on the He4 target with 200nA beam. Starting from run_num 111140 one of FCal fADC is broken. The beam was pretty stabile till ~16:37 there was some issues with RF 1L10 HV. At around 19:55 beam came back, and we continued collecting data. Then at 0:25 got call from CH, regarding DAQ rate drops dramatically from ~50kHz to ~0kHz, contacted to Sergey Furletov, and after some investigations, Sergey advised to reboot the rocTOF2, after power cycling it the rate increased to ~12kHz, but still way bellow the nominal rates (50kHz), it turns out one of FCal outer ring fADC is broken (https://logbooks.jlab.org/entry/4065801), which is causing the rate drop, Sergey excluded it from FCal trigger, which brought back the DAQ rate to nominal value. 
During this time tried to contact to expert, to schedule a rapid access to hall to replace the broken fADC. NO LUCK to reach to expert. 
Will continue running with this condition until the fADC will be replaced.
The other issue was related to gluon-03 machine, it was unresponsive from CH, but with remote connection it was OK, at around 9:38am after consulting with Sergey, rebooted the machine, now it is running as normal.

!!Stay Tuned!! 


=Saturday, October 22=
=== Past 24 hours ===
* past 24 hours we collected ~3.43B total triggers on the He4 target with 200nA beam.
* Starting from run_num 111140 one of FCal fADC is broken
* Beam was OFF from 16:37 till 19:58 issues with RF 1L10 HV
* After Beam is back continued taking data at 200nA FULL target
* ~00:15 DAQ rate dropped to 0kHz, Contacted to Sergey, 1st power cycled rocTOF2, which brings trigger rate to 12.5kHz, but still way far from nominal 50kHz 
* It turns out one of FCAL outer ring fADC is broken (trying to reach expert to replace it), will continue running with this condition until it's been replaced


=== Shift Summarys ===
* [https://logbooks.jlab.org/entry/4064773 Early Riser Oct-21]
* [https://logbooks.jlab.org/entry/4065074 Day Oct-21]
* [https://logbooks.jlab.org/entry/4065456 Swing Oct-21]
* [https://logbooks.jlab.org/entry/4065653 Owl Oct-21/22]
* [https://logbooks.jlab.org/entry/4065892 Early Riser Oct-22]


=== Issues ===
* https://logbooks.jlab.org/entry/4065801 ->FCAL fADC broken
* https://logbooks.jlab.org/entry/4066109 ->Gloun-03 reboot

=== Plan ===
*Continue taking production data at 200nA with FULL LHe target over the weekend.
      

---

This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4066146
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/halld/attachments/20221022/a06db74a/attachment.html>


More information about the Halld mailing list